OpenStudy (anonymous):

Can someone help : x^2 +y^2 + 6x -6y -46=0 graph the circle I am getting messed up when I have to complete the square and put in standard form. Thanks!

OpenStudy (anonymous):

complete the square twice. Once with x and once with y.

OpenStudy (anonymous):

When completing the square for x, ignore all the terms with y in it.

OpenStudy (anonymous):

\[x^2+6x-46=0\] Complete the square for this.

OpenStudy (anonymous):

(x^2 +6x) (6y +y^2)=46

OpenStudy (anonymous):

x(x +6)^2 = 46

OpenStudy (anonymous):

That's not how you complete the square.

OpenStudy (anonymous):

\[x^2+bx+c=0\] \[(x+\frac{b}{2})^{2}=-c+(\frac{b}{2})^2\]

OpenStudy (anonymous):

\[x^2 +y^2 + 6x -6y -46=0\] \[x^2+6x+y^2-6y=46\] \[(x+3)^2+y^2-6y=46+9\] \[(x+3)^2+(y-3)^2=46+9+9\] \[(x+3)^2+(y-3)^2=64\]

OpenStudy (anonymous):

the equation is, x^2+y^2+6x-6y-46=0 => x^2+6x+9+y^2-6x+9-18-46=0 =>(x+3)^2+(y-3)^2=18+46=64=8^2 which is of the form (x-a)^2+(y-b)^2=r^2 so center of the circle is (a,b) with radius r. here center is (-3,3) and radius is 8 unit